A car can be braked to a stop from the autobahn-like speed of 217 km/h in 170 m. Assuming the acceleration is constant, find its magnitude in (a) SI units and (b) in terms of g(c) How much time Tb is required for the braking? Your reaction time Tr is the time you require to perceive an emergency, move your foot to the brake, and begin the braking. If Tr = 400 ms, then (d) what is the ratio Tb/Tr?
Dark sunglasses delay the visual signals sent from the eyes to the visual cortex in the brain, increasing Tr(e) In the extreme case in which Tr is increased by 100 ms, how much farther along the road does the car travel during your reaction time?

Grade:12

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

V^2 = U^2 – 2aS

V=0
U= 217*5/18 m/s

S=170 m

you can find a


Kindly ask only one question at a time